A company is expected to issue dividends of $10 for the next 3 years, and 5% annual dividend growth for the 4th year and after. What is the dividend for the 5th, respectively?

A. $10.00

B. $10.50

C. 9.50

D. 11.03
Answer #1

Solution :

As per the information given in the question we have

D1 = $ 10    ;   D2 = $ 10    ;   D3 = $ 10

The dividends increase at a rate of 5 % per year after Year 3.

Calculation of Expected Dividend per share Year 4 :

As per the information given in the question we have           

D3 = $ 10

Dividend growth rate Year 4 = 5 % = 0.05

We know that D4 = D3 * ( 1 + g )

Applying the available information in the above formula we have

= $ 10 * ( 1 + 0.05 )

= $ 10 * 1.05 = $ 10.50

Thus the Expected Dividend per share for Year 4 = D4 = $ 10.50

Calculation of Expected Dividend per share Year 5 :

As per the information given in the question we have           

D4 = $ 10.50

Dividend growth rate Year 5 = 5 % = 0.05

We know that D5 = D4 * ( 1 + g )

Applying the available information in the above formula we have

= $ 10.50 * ( 1 + 0.05 )

= $ 10.50 * 1.05 = $ 11.02500

= $ 11.03 ( When rounded off to two decimal places )

Thus the Expected Dividend per share for Year 5 = D5 = $ 11.03

The solution is Option D. $ 11.03